# COMI-LINGUA: Expert Annotated Large-Scale Dataset for Multitask NLP in Hindi-English Code-Mixing

### v2

@@ Abstract
We introduce COMI-LINGUA, the largest manually annotated Hindi-English code-mixed dataset, comprising 125K+ high-quality instances across five core NLP tasks: Matrix Language Identification, Token-level Language Identification, POS Tagging, Named Entity Recognition (NER), and Machine Translation. Each instance is annotated by three bilingual annotators, yielding over 376K expert annotations with strong inter-annotator agreement (Fleiss’ Kappa 0.81). The rigorously preprocessed and filtered dataset covers both Devanagari and Roman scripts and spans diverse domains, ensuring real-world linguistic coverage. Evaluation reveals that closed-source LLMs significantly outperform traditional tools and open-source models. Notably, one-shot prompting consistently boosts performance across tasks, especially in structure-sensitive predictions like POS and NER, highlighting the effectiveness of prompt-based adaptation in code-mixed, low-resource settings. COMI-LINGUA is publicly available at this URL111 https://anonymous.4open.science/r/CodeMixing/.
@@
Rajvee Sheth, Himanshu Beniwal, Mayank Singh LINGO, Indian Institute of Technology Gandhinagar, India Correspondence: lingo@iitgn.ac.in
@@ 1 Introduction
Despite growing interest, current Hinglish datasets have critical limitations: (1) a predominant focus on Roman script, ignoring natural script variation (Begum et al., 2016; Bali et al., 2014; Srivastava et al., 2020), (2) limited scale and coverage (Srivastava and Singh, 2021a; Kumar et al., 2018; Tiwari et al., 2024; Kartik et al., 2024), (3) insufficient task diversity within single datasets (Aguilar et al., 2020; Khanujaa et al., 2020; Bohra et al., 2018; Khanuja et al., 2020), and (4) reliance on synthetic data generation rather than human annotation (Chatterjee et al., 2022; Srivastava and Singh, 2021c; Kartik et al., 2024; Sravani and Mamidi, 2023).
@@
Curation of the largest publicly available Hinglish dataset (376K manually annotated instances), released under a CC-BY-4.0 license, capturing real-world code-mixing behavior across both Roman and Devanagari scripts. Each instance is annotated by three annotators across five key NLP tasks: matrix language identification, token-level language identification, part-of-speech tagging, named entity recognition, and machine translation.
Robust benchmarking of state-of-the-art multilingual LLMs (mLLMs), including both open-weight and closed-source models—alongside traditional NLP tools, under two inference paradigms: zero-shot and one-shot in-context learning.
@@ 3.1 Raw Dataset Curation
We curated raw data from publicly accessible and licensed platforms spanning diverse categories such as news, politics, entertainment, social events, sports, and informational content, with a focus on the Indian subcontinent. Sources included prominent news portals and official digital archives, detailed in Appendix 17. The collected content was cleaned using regex-based preprocessing to remove noise such as advertisements, HTML tags, and footers, and then segmented into individual sentences. A Code-Mixing Index (CMI, Das and Gambäck (2014)) was computed for each sentence, and only those sentences with a CMI score 9 were retained to ensure a substantial degree of code-mixing. Given the under-representation of mixed Devanagari-Roman script samples in existing datasets, we also collected supplementary data to enhance coverage and linguistic diversity. This includes enriching the dataset by incorporating additional Hinglish code-mixed samples from prior works (Srivastava and Singh, 2020; Gupta et al., 2023; Singh et al., 2018c) and from Hugging-Face222https://huggingface.co/datasets/pardeep/youtube-vidoes-transcripts-hindi-english/.
@@ 3.2 Dataset Processing
At the token level, additional preprocessing steps were applied. Sentences with fewer than five tokens were discarded to eliminate non-informative content such as fragments, abbreviations, emojis, and filler phrases—commonly arising from typing errors or social media discourse. Examples of such removed content include: #GuessTheSong’, during dinner’, and ‘@enlightenedme bas ek hi’. Further data refinement was conducted during the manual annotation process (see Section 3.4 for more details).
@@ 3.3 Data Annotation
We selected five diverse annotation tasks, balancing well-established tasks with high reliability and underexplored challenges. Annotators followed detailed guidelines with examples to ensure consistency and clarity across tasks (Appendix A.3, Figure 1). The tasks are:
@@
Named Entity Recognition (NER): In the NER task, each token in a sentence is annotated with a label from a predefined set of entity types outlined in Table 1. These include conventional categories, such as Person, Location, Organization, Date/Time, and GPE (Geo-Political Entities), as well as social media-specific types like Hashtags, Mentions, and emojis. An instance of annotated entities across different types is shown in Figure 1. This allows the annotation schema to comprehensively capture the diversity and informality often observed in code-mixed social media text.
@@
Machine Translation (MT): This task involves constructing parallel translations for code-mixed sentences into three distinct formats: (i) Romanized Hindi, (ii) Devanagari-script Hindi, and (iii) standard English. The goal is to facilitate a multilingual Hinglish sentence to align with its respective translations across scripts and languages. A representative translation instance across the three formats is shown in Figure 1. Initial translation predictions were generated using the LLaMA 3.3 language model777https://github.com/meta-llama/llama-models/blob/main/models/llama3_3/MODEL_CARD.md.
@@ 3.4 Manual Data Refinement
During the annotation phase, the dataset underwent iterative refinement to ensure quality and consistency, guided by annotator feedback on instances to be excluded (see Table 6 in Appendix A.2). Sentences were removed if they (i) were monolingual English or Hindi, (ii) lacked relevant linguistic tags or named entities, contained no meaningful content, or merged multiple instances into one, or (iii) included languages other than Hindi and English, which were beyond the scope of this study. This refinement process was crucial for preserving corpus integrity and ensuring that the final dataset consisted solely of high-quality Hinglish code-mixed text. The Raw and Filtered columns in Table 3 represent the number of original instances provided for initial annotation and the final number of instances retained after annotation, respectively. The difference between these values corresponds to instances flagged by annotators as not satisfying the manual annotation criteria.
@@ 3.5 Annotation Efforts and Quality
Total / Avg. | 140,785 | 125,613 | 0.863 | 18.96 |
Model/Library | LID | POS | NER | MLI | MT | ||||||||||
claude-3.5-sonnet | 92.8 | 92.4 | 92.1 | 75.3 | 64.8 | 69.0 | 59.1 | 55.1 | 56.7 | 98.8 | 83.5 | 90.0 | 48.0 | 48.6 | 56.0 |
gpt-4o | 92.8 | 92.8 | 92.7 | 76.1 | 66.0 | 70.1 | 60.5 | 60.1 | 60.1 | 98.4 | 97.9 | 98.1 | 28.8 | 27.4 | 31.9 |
gemini-1.5-Flash | 82.9 | 40.4 | 47.9 | 73.4 | 62.4 | 66.5 | 44.2 | 44.2 | 43.8 | 98.8 | 21.4 | 33.7 | 48.1 | 28.9 | 56.9 |
LLaMA-3.3-instruct | 73.4 | 73.7 | 73.3 | 74.3 | 65.5 | 68.9 | 67.5 | 67.3 | 66.8 | 98.8 | 59.0 | 73.1 | 55.4 | 50.4 | 59.8 |
mistral-instruct | 54.5 | 39.0 | 42.4 | 10.2 | 6.72 | 7.78 | 65.1 | 41.5 | 50.2 | 98.1 | 58.7 | 72.3 | 23.5 | 5.36 | 18.05 |
command-a-03-2025 | 92.0 | 92.0 | 91.8 | 73.5 | 65.4 | 68.6 | 65.9 | 67.8 | 66.6 | 98.5 | 98.0 | 98.3 | 38.6 | 35.1 | 48.8 |
codeswitch | - | - | - | 89.1 | 87.8 | 88.2 | 81.6 | 83.1 | 81.2 | - | - | - | - | - | - |
Microsoft LID | 80.2 | 76.5 | 74.4 | - | - | - | - | - | - | - | - | - | - | - | - |
Model/Library | LID | POS | NER | MLI | MT | ||||||||||
claude-3.5-sonnet | 93.0 | 92.7 | 92.5 | 81.4 | 79.2 | 79.3 | 85.9 | 85.2 | 85.0 | 98.8 | 98.9 | 98.8 | 50.9 | 52.1 | 63.4 |
gpt-4o | 93.9 | 94.0 | 93.8 | 81.6 | 78.0 | 78.9 | 77.4 | 75.8 | 76.0 | 98.7 | 97.7 | 98.1 | 50.1 | 50.2 | 58.26 |
gemini-1.5-Flash | 80.2 | 76.5 | 74.4 | 72.9 | 64.6 | 68.0 | 66.5 | 67.5 | 66.0 | 98.4 | 40.4 | 56.4 | 44.4 | 42.9 | 66.1 |
LLaMA-3.3-instruct | 90.3 | 89.6 | 89.3 | 85.1 | 84.0 | 84.1 | 79.0 | 79.1 | 78.4 | 98.8 | 97.8 | 98.2 | 62.2 | 54.4 | 60.6 |
mistral-instruct | 72.1 | 70.0 | 70.1 | 77.3 | 66.9 | 69.8 | 65.5 | 44.4 | 52.6 | 98.3 | 88.1 | 92.7 | 30.0 | 19.5 | 18.5 |
command-a-03-2025 | 92.1 | 91.7 | 91.3 | 74.5 | 65.7 | 69.5 | 76.7 | 78.9 | 77.3 | 98.9 | 98.7 | 98.3 | 52.9 | 42.1 | 56.0 |
@@
The COMI-LINGUA consists of 125,613 high-quality instances spanning five tasks, each independently annotated by three expert annotators, yielding a total of 376,839 annotations (see Table 7) . To our knowledge, it is the largest manually annotated code-mixed dataset to date. For each task, we provide two random splits: a test set of 5,000 instances and a training set comprising the remainder.
@@ 4.1 Baseline Tools and LLMs
We conduct a comprehensive evaluation of existing tools and language models on the COMI-LINGUA Benchmark. Our experimental setup spans traditional NLP toolkits, state-of-the-art open-source LLMs, and proprietary commercial models. These systems are evaluated on their performance across five diverse Hinglish code-mixed NLP tasks, detailed in Section 3.3.
The traditional tools evaluated in this study include the Microsoft LID888https://github.com/microsoft/LID-tool designed for token-level language identification in multilingual text, and the codeswitch toolkit999https://github.com/sagorbrur/codeswitch, which provides a rule-based pipeline for annotating syntactic and semantic information in code-switched corpora. The open-source LLMs considered in our evaluation include mistral-instruct (7B) (Jiang et al., 2023) and llama-3.3-instruct (70B) (Touvron et al., 2023). In addition, we assess three commercial state-of-the-art systems: GPT-4o (Achiam et al., 2023), command-a-03-2025 (111B) (Cohere et al., 2025), Claude-3.5-Sonnet Anthropic (2024) and Gemini-1.5-Flash (Anil et al., 2023).
We create specific prompt templates for each task to generate accurate, task-aligned responses from LLMs. The prompt template includes a high-level description of the task, specific annotation or tagging rules, and illustrative examples where applicable. For each of our five tasks, we developed two prompt variants: a zero-shot version providing only task instructions and a one-shot version that includes a single demonstrative example with instructions. The prompts are presented as a system-level instruction, followed by the user-supplied test input (i.e., a code-mixed sentence or token sequence). The complete prompt template used for one task is detailed in Appendix B.

### v3

@@ Abstract
We introduce COMI-LINGUA, the largest manually annotated Hindi-English code-mixed dataset, comprising 125K+ high-quality instances across five core NLP tasks: Token-level Language Identification, Matrix Language Identification, Named Entity Recognition, Part-Of-Speech Tagging and Machine Translation. Each instance is annotated by three bilingual annotators, yielding over 376K expert annotations with strong inter-annotator agreement (Fleiss’ Kappa 0.81). The rigorously preprocessed and filtered dataset covers both Devanagari and Roman scripts and spans diverse domains, ensuring real-world linguistic coverage. Evaluation reveals that closed-weight LLMs significantly outperform traditional tools and open-weight models in zero-shot settings. Notably, one-shot prompting consistently boosts performance across tasks, especially in structure-sensitive predictions like POS and NER. Fine-tuning open-weight LLMs on COMI-LINGUA demonstrates substantial improvements, achieving up to 95.25 F1 in NER, 98.77 F1 in MLI, and competitive MT performance, setting new benchmarks for Hinglish code-mixed text. COMI-LINGUA is publicly available at this URL111https://huggingface.co/datasets/LingoIITGN/COMI-LINGUA Version 1.0, Updated till 15nd September 2025..
@@
Rajvee Sheth, Himanshu Beniwal, Mayank Singh LINGO Research Group, Indian Institute of Technology Gandhinagar, India Correspondence: lingo@iitgn.ac.in
@@ 1 Introduction
Despite growing interest, current Hinglish datasets have critical limitations: (1) a predominant focus on Roman script, ignoring natural script variation (Begum et al., 2016; Bali et al., 2014; Srivastava et al., 2020), (2) limited scale and coverage (Srivastava and Singh, 2021a; Kumar et al., 2018; Tiwari et al., 2024; Kartik et al., 2024), (3) insufficient task diversity (Aguilar et al., 2020; Khanujaa et al., 2020; Bohra et al., 2018; Khanuja et al., 2020), and (4) reliance on synthetic data generation and labeling rather than human annotation (Chatterjee et al., 2022; Srivastava and Singh, 2021c; Kartik et al., 2024; Sravani and Mamidi, 2023).
@@
Curation of the largest publicly available Hinglish dataset (376K manually annotated instances), released under a CC-BY-4.0 license, capturing real-world code-mixing behavior across both Roman and Devanagari scripts. Each instance is annotated by one annotator across one of the key NLP tasks: token-level language identification, matrix language identification, named entity recognition, part-of-speech tagging and machine translation.
Robust benchmarking of state-of-the-art multilingual LLMs (mLLMs), including both open-weight and closed-weight models—alongside traditional NLP tools, under two inference paradigms: zero-shot and one-shot in-context learning.
@@ 3.1 Raw Dataset Curation
We curated raw data from publicly accessible and licensed platforms spanning diverse categories such as news, politics, entertainment, social events, sports, and informational content, with a focus on the Indian subcontinent. Sources included prominent news portals and official digital archives, detailed in Appendix §17. The collected content was cleaned using regex-based preprocessing to remove noise such as advertisements, HTML tags, and footers, and then segmented into individual sentences. A Code-Mixing Index (CMI, Das and Gambäck (2014)) was computed for each sentence, and only those sentences with a CMI score 9 were retained to ensure a substantial degree of code-mixing. Given the under-representation of mixed Devanagari-Roman script samples in existing datasets, we also collected supplementary data to enhance coverage and linguistic diversity. This includes enriching the dataset by incorporating additional Hinglish code-mixed samples from prior works (Srivastava and Singh, 2020; Gupta et al., 2023; Singh et al., 2018c) and from Hugging-Face222https://huggingface.co/datasets/pardeep/youtube-vidoes-transcripts-hindi-english/.
@@ 3.2 Dataset Processing
At the token level, additional preprocessing steps were applied. Sentences with fewer than five tokens were discarded to eliminate non-informative content such as fragments, abbreviations, emojis, and filler phrases—commonly arising from typing errors or social media discourse. Examples of such removed content include: ‘#GuessTheSong’, ‘during dinner’, and ‘@enlightenedme bas ek hi’. Further data refinement was conducted during the manual annotation process (see Section 3.4 for more details).
@@ 3.3 Data Annotation
We selected five diverse annotation tasks, balancing well-established tasks with high reliability and underexplored challenges. Annotators followed detailed guidelines with examples to ensure consistency and clarity across tasks (Appendix §A.3, Figure 1). The tasks are:
@@
Named Entity Recognition (NER): In the NER task, each token in a sentence is annotated with a label from a predefined set of entity types outlined in Table 1. These include conventional categories, such as Person, Location, Organization, Date/Time, and GPE (Geo-Political Entities), as well as social media-specific types like Hashtags, Mentions, and emojis. An instance of annotated entities across different types is shown in Figure 1. This allows the annotation schema to comprehensively capture the diversity and informality observed in code-mixed social media text.
@@
Machine Translation (MT): This task involves constructing parallel translations for code-mixed sentences into three distinct formats: (i) Standard English, (ii) Romanized Hindi and (iii) Devanagari Hindi. The goal is to facilitate a multilingual Hinglish sentence to align with its respective translations across scripts and languages. A representative translation instance across the three formats is shown in Figure 1. Initial translation predictions were generated using the LLaMA 3.3 language model777https://github.com/meta-llama/llama-models/blob/main/models/llama3_3/MODEL_CARD.md.
@@ 3.4 Manual Data Refinement
During the annotation phase, the dataset underwent iterative refinement to ensure quality and consistency, guided by annotator feedback on instances to be excluded (see Table 8 in Appendix §A.2). We removed sentences if they (i) were monolingual English or Hindi, (ii) lacked relevant linguistic tags or named entities, contained no meaningful content, or merged multiple instances into one, or (iii) included languages other than Hindi and English, which were beyond the scope of this study. This refinement process was crucial for preserving corpus integrity and ensuring that the final dataset consisted solely of high-quality Hinglish code-mixed text. The Raw and Filtered columns in Table 3 represent the number of original instances provided for initial annotation and the final number of instances retained after annotation, respectively. The difference between these values corresponds to instances flagged by annotators as not satisfying the manual annotation criteria.
@@ 3.5 Annotation Efforts and Quality
Total / Avg. | 140,785 | 125,615 | 0.863 | 18.96 |
Model/Library | LID | MLI | NER | POS | ||||||||
Zero-shot | ||||||||||||
claude-3.5-sonnet | 92.8 | 92.4 | 92.1 | 98.8 | 83.5 | 90.0 | 59.1 | 55.1 | 56.7 | 75.3 | 64.8 | 69.0 |
gpt-4o | 92.8 | 92.8 | 92.7 | 98.4 | 97.9 | 98.1 | 60.5 | 60.1 | 60.1 | 76.1 | 66.0 | 70.1 |
gemini-1.5-Flash | 82.9 | 40.4 | 47.9 | 98.8 | 21.4 | 33.7 | 44.2 | 44.2 | 43.8 | 73.4 | 62.4 | 66.5 |
LLaMA-3.3-instruct | 73.4 | 73.7 | 73.3 | 98.8 | 59.0 | 73.1 | 67.5 | 67.3 | 66.8 | 74.3 | 65.5 | 68.9 |
mistral-instruct | 54.5 | 39.0 | 42.4 | 98.1 | 58.7 | 72.3 | 65.1 | 41.5 | 50.2 | 10.2 | 6.72 | 7.78 |
command-a-03-2025 | 92.0 | 92.0 | 91.8 | 98.5 | 98.0 | 98.3 | 65.9 | 67.8 | 66.6 | 73.5 | 65.4 | 68.6 |
codeswitch | - | - | - | - | - | - | 81.6 | 83.1 | 81.2 | 89.1 | 87.8 | 88.2 |
Microsoft LID | 80.2 | 76.5 | 74.4 | - | - | - | - | - | - | - | - | - |
One-shot | ||||||||||||
claude-3.5-sonnet | 93.0 | 92.7 | 92.5 | 98.8 | 98.9 | 98.8 | 85.9 | 85.2 | 85.0 | 81.4 | 79.2 | 79.3 |
gpt-4o | 93.9 | 94.0 | 93.8 | 98.7 | 97.7 | 98.1 | 77.4 | 75.8 | 76.0 | 81.6 | 78.0 | 78.9 |
gemini-1.5-Flash | 80.2 | 76.5 | 74.4 | 98.4 | 40.4 | 56.4 | 66.5 | 67.5 | 66.0 | 72.9 | 64.6 | 68.0 |
LLaMA-3.3-instruct | 90.3 | 89.6 | 89.3 | 98.8 | 97.8 | 98.2 | 79.0 | 79.1 | 78.4 | 85.1 | 84.0 | 84.1 |
mistral-instruct | 72.1 | 70.0 | 70.1 | 98.3 | 88.1 | 92.7 | 65.5 | 44.4 | 52.6 | 77.3 | 66.9 | 69.8 |
command-a-03-2025 | 92.1 | 91.7 | 91.3 | 98.9 | 98.7 | 98.3 | 76.7 | 78.9 | 77.3 | 74.5 | 65.7 | 69.5 |
Model | En | RH | DH | |||
Zero-shot | ||||||
claude-3.5-sonnet | 48.1 | 63.5 | 48.6 | 64.4 | 56.0 | 65.7 |
gpt-4o | 28.8 | 42.4 | 27.5 | 41.7 | 32.0 | 41.8 |
gemini-1.5-Flash | 48.2 | 66.0 | 28.9 | 50.4 | 56.9 | 66.8 |
LLaMA-3.3-instruct | 55.4 | 71.4 | 50.5 | 68.1 | 59.8 | 71.5 |
mistral-instruct | 23.5 | 49.9 | 5.4 | 25.3 | 18.1 | 40.7 |
command-a-03-2025 | 38.6 | 58.5 | 35.2 | 57.0 | 48.8 | 61.8 |
One-shot | ||||||
claude-3.5-sonnet | 50.9 | 68.5 | 52.2 | 69.4 | 63.5 | 73.4 |
gpt-4o | 50.2 | 68.6 | 50.2 | 68.4 | 58.4 | 70.1 |
gemini-1.5-Flash | 39.5 | 57.6 | 40.5 | 58.7 | 58.9 | 69.6 |
LLaMA-3.3-instruct | 62.2 | 74.8 | 54.8 | 71.3 | 60.7 | 70.3 |
mistral-instruct | 30.0 | 53.8 | 19.6 | 44.1 | 18.5 | 40.5 |
command-a-03-2025 | 52.9 | 68.6 | 42.0 | 60.4 | 56.1 | 66.6 |
Model | LID | MLI | NER | POS | ||||||||
Zero-shot | ||||||||||||
LLaMA-3.1-8B-In | 62.50 | 70.04 | 62.37 | 96.83 | 4.68 | 8.93 | 63.40 | 66.80 | 64.86 | 57.86 | 19.75 | 26.84 |
aya-expanse-8b | 51.08 | 70.55 | 59.05 | 98.71 | 59.56 | 74.25 | 54.47 | 68.27 | 59.88 | 76.92 | 29.50 | 40.55 |
Qwen2.5-7B-In | 56.43 | 69.63 | 59.13 | 98.64 | 21.20 | 34.82 | 61.02 | 65.36 | 57.37 | 68.40 | 8.70 | 9.06 |
One-shot | ||||||||||||
LLaMA-3.1-8B-In | 83.03 | 80.30 | 81.44 | 98.55 | 59.52 | 74.16 | 72.51 | 67.35 | 68.54 | 72.72 | 63.22 | 64.73 |
aya-expanse-8b | 73.03 | 71.07 | 70.48 | 98.35 | 81.36 | 89.00 | 79.73 | 81.44 | 79.18 | 55.29 | 48.70 | 48.20 |
Qwen2.5-7B-In | 57.17 | 74.85 | 64.83 | 98.49 | 61.94 | 75.74 | 74.18 | 76.30 | 74.04 | 70.46 | 59.89 | 63.09 |
Fine-tuned | ||||||||||||
LLaMA-3.1-8B-In | 95.29 | 94.57 | 94.75 | 98.07 | 88.76 | 93.05 | 95.28 | 95.24 | 95.25 | 86.66 | 86.16 | 86.17 |
aya-expanse-8b | 87.45 | 86.92 | 87.15 | 98.69 | 98.90 | 98.77 | 94.94 | 94.91 | 94.90 | 88.97 | 88.55 | 88.61 |
Qwen2.5-7B-In | 84.44 | 80.23 | 81.74 | 97.82 | 93.86 | 95.80 | 94.33 | 94.31 | 94.27 | 89.01 | 88.53 | 88.60 |
@@
The COMI-LINGUA consists of 125,615 high-quality instances spanning five tasks, each independently annotated by three expert annotators, yielding a total of 376,845 annotations (see Table 9). To our knowledge, it is the largest manually annotated code-mixed dataset to date. For each task, We provide two random splits: a test set of 5,000 instances and a training set comprising the remainder (as detailed in Table 9, Appendix §B). Zero- and one-shot prompting was evaluated only on the fixed test set, whereas fine-tuning was carried out on the training split, with performance reported on the same 5,000-instance test set.
Model | En | RH | DH | |||
Zero-shot | ||||||
LLaMA-3.1-8B-In | 38.3 | 67.5 | 15.6 | 49.2 | 7.4 | 13.5 |
aya-expanse-8b | 33.2 | 67.1 | 4.80 | 20.0 | 25.6 | 57.5 |
Qwen2.5-7B-In | 29.8 | 61.5 | 14.2 | 46.8 | 3.31 | 21.8 |
One-shot | ||||||
LLaMA-3.1-8B-In | 45.8 | 72.4 | 35.3 | 67.0 | 17.9 | 53.2 |
aya-expanse-8b | 31.7 | 65.1 | 29.7 | 63.7 | 26.4 | 59.1 |
Qwen2.5-7B-In | 30.2 | 61.7 | 18.3 | 52.8 | 35.6 | 60.7 |
Fine-tuned | ||||||
LLaMA-3.1-8BIn | 56.1 | 78.7 | 66.6 | 85.9 | 73.5 | 86.2 |
aya-expanse-8b | 55.0 | 78.1 | 62.4 | 83.7 | 69.3 | 86.0 |
Qwen2.5-7B-In | 51.9 | 76.0 | 63.9 | 84.1 | 63.8 | 78.4 |
@@ 4.1 Baseline Tools and LLMs
We conducted a comprehensive evaluation of existing tools and language models on the COMI-LINGUA Benchmark. Our experimental setup spans traditional NLP toolkits, state-of-the-art open-weight LLMs, and proprietary commercial models. These systems are evaluated on their performance across five diverse Hinglish code-mixed NLP tasks, detailed in Section 3.3.
The traditional tools evaluated in this study include the Microsoft LID888https://github.com/microsoft/LID-tool for token-level language identification and the codeswitch toolkit999https://github.com/sagorbrur/codeswitch for POS and NER tasks in multilingual text, which provides a rule-based pipeline for annotating syntactic and semantic information in code-switched corpora. The four commercial closed-weight systems considered in our evaluation include : claude-3.5-Sonnet (Anthropic, 2024), gpt-4o (Achiam et al., 2023), gemini-1.5-Flash (Anil et al., 2023) and command-a-03-2025 (111B) (Cohere et al., 2025). In addition, we assess open-weight LLMs llama-3.3-instruct (70B) (Touvron et al., 2023) and mistral-instruct (7B) (Jiang et al., 2023).
We create specific prompt templates for each task to generate accurate, task-aligned responses from LLMs. The prompt template includes a high-level description of the task, specific annotation or tagging rules, and illustrative examples wherever applicable. For each of our five tasks, we developed two prompt variants: a zero-shot version providing only task instructions and a one-shot version that includes a single demonstrative example with instructions. The prompts are presented as a system-level instruction, followed by the user-supplied test input (i.e., a code-mixed sentence or token sequence). The complete prompt template used for each task under each prompt variant is detailed in Appendix §B.